OpenStudy (anonymous):

square root 3 x square root 18

OpenStudy (anonymous):

There are multiple ways to solve this problem. Remember, \[\sqrt {n} * \sqrt {m} = \sqrt {n*m}\] There are two ways to solve this problem. You can multiply both numbers first prior to simplifying the numbers, or simplify with and then multiply.